OpenStudy (anonymous):

Suppose f^−1 is the inverse function of a differentiable function f and f(3)=7,f′(3)=7/4, then (f^−1)′(7)=?

OpenStudy (anonymous):

ok the 12 didnt work

OpenStudy (angela210793):

Wht makes u think tht the answer's 12???

OpenStudy (anonymous):

\[f'(x)\approx \frac{f(x+h)-f(x)}{x}\]

OpenStudy (angela210793):

I thought that too...but I'm not sure if that it goes this way :S

OpenStudy (anonymous):

Oh, duh - it's 3

OpenStudy (anonymous):

Lol, I'm not reading the question right.

jimthompson5910 (jim_thompson5910):

\[\Large (f^{-1})^{\prime}(x) = \frac{1}{f^{\prime}(f^{-1}(x))}\] \[\Large (f^{-1})^{\prime}(7) = \frac{1}{f^{\prime}(f^{-1}(7))}\] \[\Large (f^{-1})^{\prime}(7) = \frac{1}{f^{\prime}(3)}\] \[\Large (f^{-1})^{\prime}(7) = \frac{1}{\frac{7}{4}}\] \[\Large (f^{-1})^{\prime}(7) = \frac{4}{7}\]

OpenStudy (anonymous):

4/7 works thanks

OpenStudy (anonymous):

Props to jim_thompson5910. I learned something new today.

OpenStudy (angela210793):

me too xD

jimthompson5910 (jim_thompson5910):

Yes it's a very slick and useful formula...unfortunately the tick marks are a bit squished (next to the f), but oh well

OpenStudy (anonymous):

OH, that's why I couldn't figure out where that first line was coming from. Nice.

jimthompson5910 (jim_thompson5910):

Yeah the right side of the formula reads out to be: "1 over (f prime of f inverse of x)"

OpenStudy (anonymous):

Awesome, thanks!

jimthompson5910 (jim_thompson5910):

you're welcome